OpenStudy (anonymous):

how to solve 3x+4y=-3 y=-x-2 by using substitution method

OpenStudy (shamil98):

3x + 4y = -3 y = -x-2 so substitute the y = into 3x +4y = -3

OpenStudy (shamil98):

3x + 4(-x-2) = -3 solve the equation from there.